OpenStudy (anonymous):

multiply and simplify 4x^8(over)5p^2X25p^7(over)16x

OpenStudy (anonymous):

\[\frac{4x^8}{5p^2} \times \frac{25p^7}{16x}\] \[\frac{x^7}{4}\times 5p^5\]

OpenStudy (anonymous):

thanks your great